APACHE POINT OBSERVATORY SDSS 2.5M OBSERVING LOG Sunday December 2, 2007 (MJD 54437) ---=== OBSERVING TEAM ===--- Audrey Simmons Kaike Pan Dan Long (by phone) Steph Snedden (via phone) Frecnch Leger (via phobe) ---=== OBSERVING PLAN ===--- Science! ---=== OBSERVING SUMMARY ===--- In total about 7 hours was lost to telescope problem. We finished plates: 1873 and 2844 (the marvels plate was not declared done in the data base because the sos did not return good information, mervels team, please give us feedback on if we can declared it done), and got signal on segue faint plate 2828. Problems: 1) Very small scale changes (0.002) or focus adjustment bigger than 150 resulted losing guiding stars (see log for detail information). However, it seemed that bigger scale changes, 0.005 could be applied after a few very small ones (0.001) (exercies warmed up actuator motors?). 2) TPM message: M2 Force Limit Fault !, no related watcher message was reported. 3) cartridge manipulator had to be powered cycle 7 times tonight. 4) The west can only be closed by bypass (flopped the switch). ---=== OBSERVING LOG ===--- Afternoon: ----------- Spectro collimation set for ~0C. During afternoon checkout, we found that the west door refused to close. The door would down for about 3 to 5 inche then back up even the close control bottom was holding. Dan was called, he told us to bypass the pump, then it can be closed. Night: ------ 00:45Z cart 5, plt 2815, segue. FK5 used to find the field. Audrey did find all guiding stars. but all of them were lost after applied a scale change of 0.009. Then slew telescope back to the field and the FK 5 did not find stars. two spiralPattern did find the FK5 stars. During the second spiralPattern, we noticed that there was a message on TPM session: M2 Force Limit Fault! After the second spiralPattern, we decided to home relax the secondary. it took five attempts to get position error of actuator C down to about 450. By the time, the M2home/relax process finished, the plate expired, we switch to cart 6. 02:20Z cart 6, plate 2825, seguefaint. We could not find the guiding stars, nor FK5 star. Steph was called for her suggestions. Because in the previous M2 home/relax, the error of actuator C was a rather big, she suggested us to do another M2 home/relax. It took 9 attempted to get the position error of C actuator down to 200. After the process, we obtained a set of mirror numbers, they were looked okay. Then we slew the telescope to the field, no luck in finding stars. Then FK5 was used. We did got all guiding stars via a FK5. Sop asked for a scale change of -0.013. We did a scale change of -0.01, and lost all guiding stars. Then we send telescope to 121, 30 and obtained another set of mirror numbers which were within soectro. Then we went back to field, and found the stars, but a 0.005 scale change resulted losing guiding stars, and reset scale back to 1.0, guiding stars were back. Dan was called, and he suggested us power cycle M1 GALIL, home/relax M1. The process did not fix the problem. At the moment, We called Steph for an updating the progress, she suggested us to exit telrun and do a pair of telstop and telrun, then restarted all sessions that require tcc connections. This did not fix the problem either. After he talked to French, Dan called us back, and suggested us to power cycle PMSS control box. At the moment, cart 6 is about expire, so decided to switch to cartridge 7. 06:10Z cart 7, plt 2844, marvels. FK5 used. Most guiding stars were there, but looked very faint, and the efficiency was about 11% under photometric sky condition. The mag. information in the plugMap file were incorrect? The suitation did not improve after power cycle PMSS control box: if the focus was adjusted by more than 100, All of existing guiding stars would be lost. Any scale change would result losing guiding stars. After talking with French, Dan called us and said that they did not know for sure what was going on with the telescope, it looked like the telescope need a major surgery to back to its normal functions. Therefore, they told us if we can find something to do without change the scale, that would be great. If not, we may close for tonight, they decided to investigate the problem tomorrow because it's only 4.5 hours away from the morning twilight. No stars in guiding fiber 3, 7, and 10. The SOP GUI asked for a scale change of 0.002 to 0.003 which we could not applied without losing guiding stars. Because we could not think what else to do under the telescope condition, we decided to take some data for this plate. 1 pre-calib. 10 10 s exposures. 1 mid-calib. 5 20 s exposure (after the plate expired) 1 post-calib 08:45Z cart 4, plt 1873 (chunk 93). Stars on GTF. But required a scale change of 0.018. Some reason, we decided to give it one more try on scale change. We did by very small steps: 0.005. This time it worked!!! And the focus adjustment also behaves better than before. As soon as we started the second exposure, all guiding stars were gone, we aborted the exposure, and went to field again, no stars there, then set scale back 1.0, went to FK5, a piralPattern found the star at offset of A=3, D=10. Sticked offset, went back to the field, Sop asked a scale change of 0.015. A 0.005 scale change resulted losing all stars. Then back the scale back, and changed the scale at even small steps: 0.002 got required scale change done. then started the third exposure. The plate was DONE with three exposure, a total integration time of 3000 s. 10:35Z cart 8, plt 2828, segue faint. FK5 used. Sop asked for a scale change of -0.017. At the begining -0.002 scale change resulted losing stars. After a few excerises, -0.004 could be applied. Seeing got worse, at about 2.0". Three exposures were obtained with smallest S/N^2=13.6 on r1. 12:45Z endNight rejoiced. ---=== IMAGING RUN SUMMARY ===--- Run Time Stripe Lambda Last Flavor Comments Start End Begin End Frame ------------------------------------------------------------------------- 7207 22:23Z 22:39Z 100 O 116.35 120.24 40 ignore ---=== IMAGING RUN DETAILS ===--- ---=== SKIPPY RESULTS ===--- Run Frame nFrames stars muErr muRms nuErr nuRms rot az el --------------------------------------------------------------------------- ---=== LTMATCH RESULTS ===--- Run Field nFields alt az nGood rowMean rowSig colMean colSig rot ------------------------------------------------------------------------ ---=== SPECTROSCOPY DATA SUMMARY ===--- Summary Checked (y/n): y QA Procedures Done (y/n): y UT Exp Time flavor comment (S/N)^2 totals ========================================== b1 r1 b2 r2 ----- sequence 47621, plate -9999 ------- 22:23 47621 0.0 bias ----- sequence 47622, plate 2815 ------- 22:31 47622 10.0 flat 22:33 47623 2.0 arc ----- sequence 47624, plate -9999 ------- 23:04 47624 0.0 bias ----- sequence 47625, plate 2815 ------- 23:10 47625 0.0 bias 23:13 47626 10.0 flat 23:16 47627 2.0 arc 23:18 47628 3.0 arc 23:21 47629 3.0 arc ----- sequence 47630, plate -9999 ------- 06:46 47630 0.0 bias ----- sequence 47631, plate 2844 ------- 07:31 47631 10.0 flat 07:34 47632 2.0 arc 07:40 47633 11.0 target 07:43 47634 11.0 target 07:47 47635 11.0 target 07:50 47636 11.0 target 07:53 47637 11.0 target 07:56 47638 11.0 target 07:59 47639 11.0 target 08:02 47640 11.0 target 08:05 47641 11.0 target 08:08 47642 11.0 target 08:12 47643 10.0 flat 08:15 47644 2.0 arc 08:19 47645 21.0 target 08:23 47646 21.0 target 08:26 47647 21.0 target 08:29 47648 21.0 target 08:33 47649 21.0 target 08:36 47650 10.0 flat 08:39 47651 2.0 arc ----- sequence 47652, plate 1873 ------- 18.1 15.8 19.8 17.9 09:07 47652 10.0 flat 09:09 47653 2.0 arc 09:28 47654 901.0 target 10:09 47655 900.0 target 10:32 47656 1203.0 target ----- sequence 47657, plate 2828 ------- 14.8 13.6 19.4 18.9 11:01 47657 10.0 flat 11:04 47658 2.0 arc 11:33 47659 1502.0 target 12:00 47660 1500.0 target 12:03 47661 10.0 flat 12:06 47662 2.0 arc 12:24 47663 900.0 target 12:27 47664 10.0 flat 12:29 47665 2.0 arc ---=== TELESCOPE OFFSETS AND SCALE I ===--- Time Instrument Az Alt Rot Scale pos offset pos offset pos offset ------------------------------------------------------------------------------ 08:00Z 7 2844 229.14 -0.0125 53.28 0.0011 79.58 0.0000 1.000000 10:06Z 4 1873 170.08 -0.0147 63.89 0.0023 -15.91 0.0055 1.000140 11:11Z 8 2828 -9.52 -0.0120 52.46 -0.0006 172.00 0.0137 0.999960 ---=== TELESCOPE OFFSETS AND SCALE II ===--- ---=== FOCUS LOG ===--- setmir piston Temp Wind Time Inst scale M1 M2 Foc Az Alt (C) MPH Dir filt fwhm ------------------------------------------------------------------------------ 08:00Z 7 2844 1.00000 0 1120 -50 229 53.4 3.9 1 318 -- 1.6 10:06Z 4 1873 1.00014 -1383 -51 -50 170 63.9 1.3 2 287 -- 1.3 11:11Z 8 2828 0.99996 395 1425 -75 -10 52.5 2.5 3 28 -- 1.9 ---=== WEATHER LOG ===--- Wind Time Temp F Dewp F MPH Direction Dust DIMM Sky 22:04Z 33 21 0 339 (NNW) 1057 - clear 22:34Z 31 21 5 339 (NNW) 1325 - 23:04Z 31 22 1 41 (NE) 1363 - 23:35Z 32 12 16 27 (NNE) 1108 - 00:11Z 34 -10 18 36 (NE) 203 - 00:42Z 34 -9 14 21 (NNE) 178 - 01:13Z 36 -11 22 20 (NNE) 96 - 01:44Z 35 -11 14 32 (NNE) 108 - 02:17Z 35 -12 19 32 (NNE) 91 - 02:48Z 36 -11 16 9 (N) 209 - 03:18Z 37 -11 8 21 (NNE) 124 - 03:49Z 37 -11 10 16 (NNE) 114 - 04:19Z 37 -11 17 20 (NNE) 82 - 04:49Z 39 -8 13 35 (NE) 53 - 05:21Z 39 -8 1 314 (NW) 95 - 05:51Z 39 -10 1 128 (SE) 81 - 06:21Z 39 -9 2 183 (S) 85 - 07:12Z 34 4 2 180 (S) 425 - 07:42Z 35 -11 1 84 (E) 266 - 08:13Z 38 -8 4 272 (W) 84 - 08:45Z 34 4 4 220 (SW) 400 - 09:17Z 34 4 0 282 (WNW) 426 - 09:48Z 34 1 1 344 (NNW) 412 - 10:18Z 35 -1 5 342 (NNW) 402 - 10:48Z 36 -9 7 350 (NNW) 126 - 11:19Z 40 -7 8 30 (NNE) 93 - 11:49Z 36 -9 3 315 (NW) 142 - 12:19Z 36 -9 6 349 (NNW) 146 - 12:52Z 41 -6 13 7 (N) 76 - clear ---=== TELESCOPE STATUS ===--- 23:30Z - Began ventilation. Doors and louvers opened, fans on. 00:10Z - Enclosure off. 12:40Z - Enclosure on, doors and louvers closed, fans off. No interlocks bypassed. Status at: 12:50Z Telescope stowed at: 30,121,00 mount Instrument mounted: Cartridge 8 Counterweights at: 270 lbs LN2 autofill systems: Connected and turned on. 180L LN2 dewar scales: Spectro 83 lbs, 25 psi Imager 186 lbs, 21 psi ---=== SOFTWARE USED ===--- IOP/SOP: v4_38_0 Watcher: v2_36_0 MCP: v6_4_0 TPM: tpm_v3_5_0 AstroDa: v15_13_1 TCC: TCC 2.7.2.1 August 6 2004 sdssProcedures: SoS: v5_1_8 plate-mapper: v4_3_1 ---=== MIRROR NUMBERS ===--- PRIMARY: -------- Scale: 1.000000 MIGS TONIGHT NOMINAL Axial A -5.2150 -5.2130 Axial B -0.2135 -0.2135 Axial C -0.0015 -0.0016 Trans D 11.6450 11.6400 Lateral E 9.9695 9.9568 Lateral F 11.5443 11.5316 GALILS Commanded: 4600. -2000. 2650. -9300. -9800. -9800. Actual: 4600. -2000. 2650. -9300. -9800. -9800. SETMIR VALUES PriDesOrient: 0.00 0.00 0.00 0.00 0.00 PriOrient: 0.00 0.00 0.00 0.00 0.00 SECONDARY: ---------- Focus: 0.00 Air Temp.: 0.5 Alt.: 30.004383 MIGS TONIGHT NOMINAL Axial A 0.7976 0.7973 Axial B 0.6038 0.6041 Axial C 0.5430 0.5425 Trans D 0.5999 0.5999 GALILS Commanded: 1799453. 1310623. 1642723. -9450. -4400. Actual: 1798628. 1309538. 1643409. -9472. -4376. SETMIR VALUES SecDesOrient: 1257.00 -60.00 -140.00 -190.00 200.12 SecOrient: 1257.32 -59.42 -140.04 -189.49 199.75 ----------------- PRIMARY: -------- Scale: 1.000000 MIGS TONIGHT NOMINAL Axial A -5.2160 -5.2130 Axial B -0.2136 -0.2135 Axial C -0.0016 -0.0016 Trans D 11.6550 11.6400 Lateral E 9.9822 9.9568 Lateral F 11.5443 11.5316 GALILS Commanded: 4600. -2000. 2650. -9300. -9800. -9800. Actual: 4600. -2000. 2650. -9300. -9800. -9800. SETMIR VALUES PriDesOrient: 0.00 0.00 0.00 0.00 0.00 PriOrient: 0.00 0.00 0.00 0.00 0.00 SECONDARY: ---------- Focus: 0.00 Air Temp.: 2.7 Alt.: 75.383256 MIGS TONIGHT NOMINAL Axial A 0.7978 0.7973 Axial B 0.6063 0.6041 Axial C 0.5377 0.5425 Trans D 0.6012 0.5999 GALILS Commanded: 1807176. 1305306. 1623832. -9450. -4400. Actual: 1798628. 1309538. 1643409. -9472. -4376. SETMIR VALUES SecDesOrient: 1257.00 -60.00 -140.00 -190.00 200.12 SecOrient: 1252.96 -53.26 -145.32 -192.95 195.69 ----------- PRIMARY: -------- Scale: 1.000000 MIGS TONIGHT NOMINAL Axial A -5.2160 -5.2130 Axial B -0.2136 -0.2135 Axial C -0.0016 -0.0016 Trans D 11.6590 11.6400 Lateral E 9.9822 9.9568 Lateral F 11.5443 11.5316 GALILS Commanded: 4600. -2000. 2650. -9300. -9800. -9800. Actual: 4600. -2000. 2650. -9300. -9800. -9800. SETMIR VALUES PriDesOrient: 0.00 0.00 0.00 0.00 0.00 PriOrient: 0.00 0.00 0.00 0.00 0.00 SECONDARY: ---------- Focus: 0.00 Air Temp.: 2.4 Alt.: 70.727271 MIGS TONIGHT NOMINAL Axial A 0.7972 0.7973 Axial B 0.6068 0.6041 Axial C 0.5411 0.5425 Trans D 0.6005 0.5999 GALILS Commanded: 1801028. 1304437. 1617558. -9450. -4400. Actual: 1798628. 1309538. 1643409. -9472. -4376. SETMIR VALUES SecDesOrient: 1257.00 -60.00 -140.00 -190.00 200.12 SecOrient: 1249.44 -51.99 -143.93 -192.03 194.84 ---=== PROBLEMS IN DETAIL ===--- --------------------------------------------------------------- Spectro Collimation with Cartridge 5 Before: sp1> -6352 -6542 -7010 T= 7.6 sp2> -1827 -2743 -2907 T= 5.4 Average T= 6.5 Spectro collimation set for 0C Difference of 6.5 degrees (colder) = -6500 steps IDL> collimate Camera Steps/Degrees Focus Mean Offset MechPiston Result r1 1445 OUT -0.16 -5055 r2 3535 OUT -0.39 -2965 b1 -2.3 IN 0.06 ... b2 -7.9 OUT 0.22 ... After: sp1> -11407 -11596 -12075 sp2> -4792 -5701 -5872 sop> caget tpm_TM_PMSSAvg tpm_TM_TelairAvg.VALA tpm_TM_PMSSAvg = 4.715 tpm_TM_TelairAvg.VALA = 5.48 ---------------------------------------------------------------